A trailer of mass 250 kg is accelerated from rest by a 50 N force from the right for a period of
20 seconds. What is the final velocity of the trailer?
1
Expert's answer
2020-11-05T10:44:06-0500
"v=at=\frac{F}{m}t\\v=\frac{50}{250}20=4\frac{m}{s}"